# DWTGS: Rethinking Frequency Regularization for Sparse-view 3D Gaussian Splatting

### v1

@@ Abstract
Index Terms:
Sparse-view 3DGS, frequency domain, wavelet transform, Fourier transform, frequency regularization
@@ I Introduction
FreeNeRF [16] first demonstrates the surprising effectiveness of prior-free frequency regularization. It demonstrates that sparse‑view NeRF suffers from HF overfitting, as simply suppressing HF inputs and relying only on LF in early training markedly improves results. Building on this idea at the loss‑function level, where supervision is decoupled from the specific inputs (e.g., 3D coordinates, Gaussians), FreGS [2] and PGDGS [17] introduce fully supervised, Fourier‑space losses that split LF and HF components, injecting the HF term only after an initial LF‑focused phase. However, because learning slowly-varying LF structures is relatively simple, this supervision ends up being biased towards HF details (middle column, Fig. 1), reverting to the HF overfitting weakness identified by FreeNeRF.
@@ II Related Works
Discrete Wavelet Transform (DWT) for Neural Rendering. Recently, the DWT has gained increasing attraction in deep computer vision frameworks as it enables disentangled frequency learning [20], multi-scale understanding [21], lossless feature downsampling [22], computational efficiency [23] and flexible representation learning [24, 25]. The DWT has also been leveraged in NeRF-based frameworks, e.g., for compression [26] and fine detail enhancement [27, 28, 29]. While still limited, extensions to 3DGS are also being explored, e.g., for Gaussian sampling [30], fine detail enhancement [31] and coarse-to-fine learning [32]. Within this line of works, our DWTGS framework novelly introduces the DWT to sparse-view 3DGS at the loss function level, representing an effective, spatially-aware method of frequency regularization.
@@ III Preliminary Background
where $\mathbf{X}^{gt}$ and $\mathbf{X}$ denote the ground-truth and render images from the same training viewpoint, respectively. $\mathcal{L}_{1}$ is the MAE loss, and $\mathcal{L}_{D\text{-}SSIM}$ supervises perceptual similarity [35], with a balancing term $\lambda$. To adaptively control the Gaussians’ densities, 3DGS proposes the ADC scheme [1], which can remedy a problem termed “over-reconstruction”. This happens when high-variance 3D regions are covered only by a few coarse-sized Gaussians, leading to unseemly artifacts [2].
@@ III-B Discrete Wavelet Transform (DWT)
Figure -1: 1-level (b) and 2-level DWT subbands (c) of a cropped image region (a) (“trex” scene, LLFF [36] dataset). Each subband provides directional frequency information, and gets coarser as the DWT level increases.
@@
where $\mathbf{L}_{(\cdot)}$ and $\mathbf{H}_{(\cdot)}$ are the low-pass and high-pass matrices that filter the columns or rows of $\mathbf{X}$, respectively. The subscript $\{0,1\}$ denotes filtering along columns or rows. Each filtering matrix is constructed by shifting columns or rows of a common 1D wavelet filter [37]. The LL subband is obtained by applying low-pass filters along both rows and columns, preserving the image’s coarse structure. The LH and HL subbands result from applying a high-pass filter in one direction and a low-pass filter in the other, highlighting horizontal and vertical details, respectively. Lastly, the HH subband is obtained by high-pass filtering in both directions, capturing fine diagonal details. By repeating (2) for the resulting LL subband, the multi-level DWT can be achieved. Fig. -1 provides a visual example of the 2-level DWT subbands.
@@ IV-A Fourier-space Frequency Regularization
FreGS [2] shows that the densification scheme introduced in Section III-A is inadequate to remedy the over-reconstruction issue and represent high-frequency (HF) details, as the pixel-wise supervisions in (1) have a low-frequency (LF) bias [33]. Therefore, to disentangle frequency learning, it proposes an auxiliary loss that supervises between the ground-truth (GT) and render frequency spectra with magnitude |()||($\cdot$)|| ( ⋅ ) | and phase ($\angle$($\cdot$)∠ ( ⋅ ) terms:
@@
However, evidently, $\mathcal{L}_{FreGS}$ requires four terms of significantly different ranges (between LP and HP, then between magnitude and phase). When used alongside the main loss $\mathcal{L}_{3DGS}$, this requires extensive tuning for the weights. Furthermore, it requires defining parameters of the Fourier-space filtering operations that perform LF-HF separation. Those parameters are not detailed in the papers [2, 17], making the method’s implementation ambiguous and subjected to tuning. As seen in Fig. III-B, we implement element-wise Gaussian masks for LF-HF separation, but the mask’s standard deviation is hardly interpretable in the Fourier space.
@@ IV-B Wavelet-space Joint Spatial-Frequency Regularization
Figure 1: Visual ablations on a novel view region (a) (“trex” scene, LLFF [36] dataset). LF-centric wavelet-space losses at c), d) and e) enforce better structural consistency than Fourier-space loss at b). However, fully supervising HF at d) degrades image quality. Instead, in e), the HF is self-supervised and enforced to be sparse, which yields the best quality.
@@ V-A Datasets & Implementation Details
Datasets. DWTGS is evaluated on the LLFF [36] (3 views), Mip-NeRF 360 [43] (12 views) and Blender [3] (8 views) datasets. Novel view synthesis performance of 3DGS is evaluated on held-out test images. We adopt PSNR, SSIM [35] and LPIPS [45] as evaluation metrics, and mainly compare DWTGS with its Fourier-space predecessor, FreGS [2, 17].
Implementation Details. Our implementations are built on DropGaussian [44]. Due to lack of code availability, we re-implement the Fourier-space losses ourselves. For LF-HF separation, we leverage a LP Gaussian element-wise mask that only retains 50% of all frequencies, and use its inverse as the HP mask. For progressive HF supervision, we attenuate the HP mask based on the current training iteration, as illustrated by the donut shape in Fig. III-B. We only supervise the HF after 5K iterations. The terms $\lambda_{LP}$ and $\lambda_{HP}$ are set to 0.01 to balance with the main objective $\mathcal{L}_{3DGS}$. Those parameters are chosen after extensive experiments to ensure the best performances. In implementing the wavelet-space losses, we set the balancing term $\lambda_{LL}$ to 0.5 and use $N=2$ DWT levels. We regularize the HF with (5) by rendering novel views every 10 iterations, and stop after 5K iterations to save computation. Finally, we used the Haar wavelet [37] for the DWT.
@@ V-B Quantitative Results
Tables I and II show the quantitative results. We re-ran all methods on our machine, except 3DGS. Generally, across all sparse-view datasets, Fourier-space losses provide little performance increases, while our wavelet-space losses consistently boost performance non-trivially, by about 0.3-0.4 PSNR.

### v3

@@ I Introduction
FreeNeRF [16] first demonstrates the surprising effectiveness of prior-free frequency regularization. It demonstrates that sparse‑view NeRF suffers from HF overfitting, as simply suppressing HF inputs and relying only on LF in early training markedly improves results. Building on this idea at the loss‑function level, where supervision is decoupled from the specific inputs (e.g., 3D coordinates, Gaussians), FreGS [2] and PGDGS [17] utilize fully supervised, Fourier‑space losses that split LF and HF components, injecting the HF term only after an initial LF‑focused phase. However, because learning slowly-varying LF structures is relatively simple, this supervision ends up being biased towards HF details (middle column, Fig. 1), reverting to the HF overfitting weakness identified by FreeNeRF.
@@ II Related Works
Discrete Wavelet Transform (DWT) for Neural Rendering. Recently, the DWT has gained increasing attraction in deep computer vision frameworks as it enables disentangled frequency learning [20], multi-scale understanding [21] and flexible representation learning [22, 23]. While still limited, extensions of the concept to 3DGS are also being explored, e.g., for fine detail enhancement [24, 25] and coarse-to-fine learning [26]. Within this line of works, our DWTGS framework novelly introduces the DWT to sparse-view 3DGS at the loss function level, representing an effective, spatially-aware method of frequency regularization.
@@ III Preliminary Background
where $\mathbf{X}^{gt}$ and $\mathbf{X}$ denote the ground-truth and render images from the same training viewpoint, respectively. $\mathcal{L}_{1}$ is the MAE loss, and $\mathcal{L}_{D\text{-}SSIM}$ supervises perceptual similarity [29], with a balancing term $\lambda$. To adaptively control the Gaussians’ densities, 3DGS proposes the ADC scheme [1], which can remedy “over-reconstruction”, which happens when high-variance 3D regions are covered only by a few coarse-sized Gaussians, leading to unseemly artifacts [2].
@@ III-B Discrete Wavelet Transform (DWT)
Figure -1: 1-level (b) and 2-level DWT subbands (c) of an image region (a) [30]. Each subband provides directional frequency information, and gets coarser with increasing DWT level.
@@
where $\mathbf{L}_{(\cdot)}$ and $\mathbf{H}_{(\cdot)}$ are the low-pass and high-pass matrices that filter the columns or rows of $\mathbf{X}$, respectively. The subscript $\{0,1\}$ denotes filtering along columns or rows. Each filtering matrix is constructed by shifting columns or rows of a common 1D wavelet filter [31]. The LL subband is obtained by applying low-pass filters along both rows and columns, preserving the image’s coarse structure. The LH and HL subbands result from applying a high-pass filter in one direction and a low-pass filter in the other, highlighting horizontal and vertical details, respectively. Lastly, the HH subband is obtained by high-pass filtering in both directions, capturing fine diagonal details. By repeating (2) for the resulting LL subband, the multi-level DWT can be achieved, as shown in Fig. -1.
@@ IV-A Fourier-space Frequency Regularization
FreGS [2] shows that the densification scheme introduced in Section III-A is inadequate to remedy the over-reconstruction issue and represent high-frequency (HF) details, as the pixel-wise supervisions in (1) have a low-frequency (LF) bias [27]. Therefore, to disentangle frequency learning, it proposes an auxiliary loss that supervises between the ground-truth (GT) and render frequency spectra with magnitude $|(\cdot)|$ and phase $\angle(\cdot)$ terms:
@@
However, evidently, Fourier-based losses like $\mathcal{L}_{FreGS}$ require four terms of significantly different ranges (between LP and HP, then between magnitude and phase). When used alongside the main loss $\mathcal{L}_{3DGS}$, this requires extensive tuning for the weights. Furthermore, it requires defining parameters of the Fourier-space filtering operations that perform LF-HF separation. Those parameters are not detailed in the papers [2, 17], making the method’s implementation ambiguous and subjected to tuning. As seen in Fig. III-B, we implement element-wise Gaussian masks for LF-HF separation, but the mask’s standard deviation is hardly interpretable in the Fourier space.
@@ IV-B Wavelet-space Joint Spatial-Frequency Regularization
Figure 1: Visual ablations on a novel view region (a) [30]. LF-centric wavelet-space losses at c), d) and e) enforce better structural consistency than Fourier-space loss at b). However, fully supervising HF at d) degrades image quality. Instead, in e), the HF is self-supervised and enforced to be sparse, which yields the best quality.
@@ V-A Datasets & Implementation Details
Datasets. DWTGS is evaluated on the LLFF [30] (3 views), Mip-NeRF 360 [38] (12 views) and Blender [3] (8 views) datasets. Novel view synthesis performance of 3DGS is evaluated on held-out test images. We adopt PSNR, SSIM [29] and LPIPS [40] as evaluation metrics, and mainly compare DWTGS with its Fourier-space predecessor, FreGS [2].
Implementation Details. Our implementations are built on DropGaussian [39], which is based on the original 3DGS [1], but with a Gaussian random dropping strategy that boosts sparse-view performances. Due to lack of code availability, we re-implement the Fourier-space losses ourselves. For LF-HF separation, we leverage a LP Gaussian element-wise mask that only retains 50% of all frequencies, and use its inverse as the HP mask. For progressive HF supervision, we attenuate the HP mask based on the current training iteration, as illustrated by the donut shape in Fig. III-B. We only supervise the HF after 5K iterations. The terms $\lambda_{LP}$ and $\lambda_{HP}$ are set to 0.01 to balance with the main objective $\mathcal{L}_{3DGS}$. Those parameters are chosen after extensive experiments to ensure the best performances. In implementing the wavelet-space losses, we set $\lambda_{LL}$ to 0.5 and use $N=2$ DWT levels. We regularize the HF with (5) by rendering novel views every 10 iterations, and stop after 5K iterations to save computation. Finally, we used the Haar wavelet [31] for the DWT.
@@ V-B Quantitative Results
Tables I and II show the quantitative results, averaged across all scenes of each dataset. We re-ran all methods on our machine, except 3DGS. Generally, across all datasets, Fourier-space losses provide little performance increases, while our wavelet-space losses increase the PSNR by 0.3-0.4.
@@ VI Conclusion
Acknowledgements. The first author of this work was financially supported by the Vingroup Science and Technology Scholarship Program for Overseas Study for Master’s and Doctoral Degrees.